OpenStudy (anonymous):

marvin invested $6000 more at 5% than he invested at 8%. his total income was $2420 . how much did he investe at each rate? i know that Interest = principal * rate* time. is this question solvable, since time is not given? if t = 1, then its solvable, and i got the the amount invested at 8% is 16307.76

OpenStudy (anonymous):

i get x = amount invested at 5 y = amount invested at 8% .05 ( 6000 + y) + .08y = 2420

OpenStudy (anonymous):

but dont we need to take time into account here

OpenStudy (anonymous):

the interest accumulated on left = total interest on right

OpenStudy (anonymous):

so it should be .05 ( 6000 + y)*t + .08y*t = 2420 , where 2420 has the time built in

OpenStudy (anonymous):

.05 ( 6000 + y) + .08y = 2420 thats correct

OpenStudy (anonymous):

they really should tell you the time period, but since its not given you just assume a year ( or however long it takes to get 8% interest )

OpenStudy (anonymous):

He invested P dollars at 8% for t years and got a return of 0.08Pt dollars. Then he invested P + 6000 dollars at 5% interest for t years and got a return of (0.05P + 300)t dollars. So his total income is (0.05P + 300)t + 0.08Pt = 2420 Without knowing how long the investment lasted, we can't tell the amount he originally invested.

OpenStudy (anonymous):

elecengineer, 8 percent is given initially , so

OpenStudy (anonymous):

what do you mean, how long it takes to get 8 percent interest

OpenStudy (anonymous):

just one time period

OpenStudy (anonymous):

i think the interest rate is stipulated

OpenStudy (anonymous):

if the answer says he earns 8% then that means it was in the bank for long enough to earn 8% , if that makes sense.

OpenStudy (anonymous):

no all they are saying is that there is a certain period of time , over which one investment gets 8% and the other 5%

OpenStudy (anonymous):

actually time is not really mentioned

OpenStudy (anonymous):

but i guess in these types of problems, we just assume time is year, since it isnt given

OpenStudy (anonymous):

which is why you assume , they wouldnt mark it down it wasnt specified , but you dont actually need to the know actual time to answer the question anyway

OpenStudy (anonymous):

because both investments are for the same time, it has no effect

OpenStudy (anonymous):

no, read what he said, james

OpenStudy (anonymous):

He invested P dollars at 8% for t years and got a return of 0.08Pt dollars. Then he invested P + 6000 dollars at 5% interest for t years and got a return of (0.05P + 300)t dollars. So his total income is (0.05P + 300)t + 0.08Pt = 2420 Without knowing how long the investment lasted, we can't tell the amount he originally invested.

OpenStudy (anonymous):

james i agree, two unknowns , cant solve

OpenStudy (anonymous):

elecengineer, i think the time does have an effect.

OpenStudy (anonymous):

these problems though are ubiqutous in baby algebra books. a system of equations, something in terms of rate and total. except we get this problem in this case

OpenStudy (anonymous):

i say baby algebra, to distinguish from algebra (group theory , rings, fields, )

OpenStudy (anonymous):

you can find the original amount invested P in terms of t (which is in years): P = (2420 - 300t)/0.13t and hence get an upper bound on t: t <= 8.0667 years

OpenStudy (anonymous):

how>?

OpenStudy (anonymous):

because he invested a non-negative amount of money

OpenStudy (anonymous):

oh , so youre looking at that function

OpenStudy (anonymous):

y = (2420 - 300x)/(0.13x)

OpenStudy (anonymous):

x>0 and y> 0

OpenStudy (anonymous):

wait, P goes to infinity as t approaches zero